Percent of Change:
What number increased by 30% equals 260? I know that the answer is 200 but how can I show my work?

Answer by psbhowmick(878)

Let the number = x.
30% of x = 0.3x
If the number is increased by 30% then it becomes = x + 0.3x = 1.3x.
This number = 260.
So 1.3x = 260
or x = 260/1.3 = 200

This is the way to show the workings.